Val wanted to know the average shearing strength, in pounds (lbs.), of a particular kind of rivet sold in a hardware store. He tested 20 rivets as samples and got the following results.




518




490




513




598




510




532




512




455




500




512




501




487




498




496




500




498




515




520




497




502





Construct a confidence interval for the population mean using 99% confidence.

1
Expert's answer
2022-05-06T12:56:06-0400
"\bar{x}=\dfrac{1}{20}(
518+490+513+598+510"

"+532+512+455+500+512+501"

"+487+498+496+500+498+515"

"+520+497+502)=507.7"

"s^2=\dfrac{1}{20-1}\displaystyle\sum_{i=1}^{20}(x_i-507.7)^2=\dfrac{13280.2}{19}"

"s=\sqrt{\dfrac{13280.2}{19}}\approx26.4378"

The critical value for "\alpha = 0.01" and "df = n-1 = 19" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 2.860935."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(507.7-2.860935\times\dfrac{26.4378}{\sqrt{20}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"507.7+2.860935\times\dfrac{26.4378}{\sqrt{20}})"

"=(490.7871, 524.6129)"

Therefore, based on the data provided, the 99% confidence interval for the population mean is "490.7871< \mu < 524.6129," which indicates that we are 99% confident that the true population mean "\mu" is contained by the interval "(490.7871, 524.6129)."